i have $2.05 in quarters and dimes. i have 4 more quarters than dimes. what is the equation for this? trying to help my niece , it's truly been years.thanks
Found 2 solutions by nerdybill, scott8148:
Answer by nerdybill(7384)

i have $2.05 in quarters and dimes. i have 4 more quarters than dimes. what is the equation for this?
.
Let d = number of dimes
then from "i have 4 more quarters than dimes" we have
d+4 = number of quarters
.
Our equation comes from: "i have $2.05 in quarters and dimes":
.10d + .25(d+4) = 2.05
.10d + .25d+1 = 2.05
.35d+1 = 2.05
.35d = 1.05
d = 1.05/.35
d = 3 (number of dimes)
.
Quarters:
d+4 = 3+4 = 7

Answer by scott8148(6628)

q = d + 4

10d + 25q = 205

substituting ___ 10d + 25(d+4) = 205 ___ 10d + 25d + 100 = 205 ___ 35d = 105 ___d = 3

substituting ___ q = (3) + 4 = 7